Tony is 8 years older than James. In ten years time he will be 3 times James' current age. How old is tony now? Please help ASAP because I have an exam on Monday. Thanks
Answer by richwmiller(17219)

t=j+8
t+10=3j
j+8+10=3j
j+18=3j
18=2j
j=9
t=17
check
17+10=3*9
27=27
ok